Will the following reaction occur spontaneously at 25°C, given that [Fe2+] = 0.60 M and [Cd2+] = 0.010 M? Cd(s) + Fe2+(aq) → Cd2+(aq) + Fe(s)

1
Expert's answer
2021-12-21T21:07:02-0500

Cd(s) + Fe2+(aq) → Cd2+(aq) + Fe(s)

Oxidation: Cd → Cd2+ + 2e-

Reduction: 2e- + Fe2+ → 2Fe

E0 = E0Fe2+/Fe - E0Cd2+/Cd = -0.44 - (-0.40) = -0.04 V

E = E0 - 0.0257 V / n x lnQ

E = -0.04 V - 0.0257 V / 2 x ln(0.010/0.60) = 0.013 V

E > 0 (spontaneous)